Stock A has a beta of 1.70 and an expected return of 19.5 percent. Stock B has a beta of 1.10 and an expected return of 14 percent. If CAPM holds, what should the return on the market and the risk-free rate be?
Answer #1

Using CAPM Model,

0.195 = Rf + 1.70(Market Risk Premium)

0.14 = Rf + 1.10(Market Risk Premium)

Equating both,

0.195 = 0.14 + (1.70 - 1.10)(Market Risk Premium)

Market Risk Premium = 0.0917

0.195 = Rf + 1.70(0.0917)

Rf = 3.91%

Risk Free Rate = 3.91%

Market Rate = 0.0917 + 0.0391

Market Rate = 13.08%